# The 4D Human Embryonic Brain Atlas: spatiotemporal atlas generation for rapid anatomical changes using first-trimester ultrasound from the Rotterdam Periconceptional Cohort

### v1

@@ Abstract
Early brain development is crucial for lifelong neurodevelopmental health. However, current clinical practice offers limited knowledge of normal embryonic brain anatomy on ultrasound, despite the brain undergoing rapid changes within the time-span of days. To provide detailed insights into normal brain development and identify deviations, we created the 4D Human Embryonic Brain Atlas using a deep learning-based approach for groupwise registration and spatiotemporal atlas generation. Our method introduced a time-dependent initial atlas and penalized deviations from it, ensuring age-specific anatomy was maintained throughout rapid development. The atlas was generated and validated using 831 3D ultrasound images from 402 subjects in the Rotterdam Periconceptional Cohort, acquired between gestational weeks 8 and 12. We evaluated the effectiveness of our approach with an ablation study, which demonstrated that incorporating a time-dependent initial atlas and penalization produced anatomically accurate results. In contrast, omitting these adaptations led to anatomically incorrect atlas. Visual comparisons with an existing ex-vivo embryo atlas further confirmed the anatomical accuracy of our atlas. In conclusion, the proposed method successfully captures the rapid anotomical development of the embryonic brain. The resulting 4D Human Embryonic Brain Atlas provides a unique insights into this crucial early life period and holds the potential for improving the detection, prevention, and treatment of prenatal neurodevelopmental disorders.
Keywords spatiotemporal atlas  $\cdot$ embryonic brain  $\cdot$ brain development  $\cdot$ groupwise image registration  $\cdot$ first-trimester ultrasound  $\cdot$ deep learning
@@ 1 Introduction
We extended the work by Dalca et al. (2019) by introducing a time-dependent initial atlas and penalized deviations from this initial atlas. This is done to ensure that the rapid anatomical changes that happen in a time-span of days are maintained in the atlas. The main contributions of research presented in this paper are:
@@
    a deep learning-based approach for spatiotemporal atlas creation adapted to account for rapid anotomical changes [^1] Preliminary results were published in Bastiaansen et al. (2022b). There, we build upon the work by Dalca et al. (2019) but instead of jointly registering images to the atlas and generating the atlas, the images were registered to the atlas, and the atlas was updated iteratively.;
@@
    the 4D Human Embryonic Brain Atlas, the first spatiotemporal brain atlas covering the first trimester based on ultrasound imaging;
@@
    an ablation study evaluating the effect of the time-dependent initial atlas;
@@
    an evaluation of the anatomical correctness of the resulting atlas by comparing its volume to a reference head volume (HV) curve established in previous research (Koning et al., 2016) and visual comparison to the 3D Embryo Atlas by de Bakker et al. (2016) and to the Fetal Brain Atlas by Namburete et al. (2023);
  5. 5.
    a study to asses whether the atlas can capture known differences in morphology in the case of maternal overweight and obesity during early pregnancy using Voxel-Based Morphometry (Ashburner and Friston, 2000).
@@ 2.1 Atlas $A_{t}$ and images $I_{i,t}$
Figure 1: Overview of the proposed method to generate a spatiotemporal atlas. The first network is the atlas generator network, which takes as input the time $t$ in days and the initial atlas $A_{t}^{0}$ and outputs the atlas $A_{t}=A_{t}^{0}+A_{t}^{g}$. The second network is the nonrigid registration network, which takes as input the atlas $A_{t}$ and an image $I_{i,t}$ for subject $i$ imaged at time $t$ and gives as output the deformation field $\phi_{i,t}$ and its inverse.
@@ 2.2 The time-dependent initial atlas $A_{t}^{0}$
The network generates the atlas by learning the deviation $A_{t}^{g}$ from the initial atlas, which is an input of the network. We create a time-dependent initial atlas, instead of using a single initial atlas as proposed by Dalca et al. (2019), to account for the rapid morphological changes (e.g., the appearance of the choroid plexus, change of curvature of the embryonic body). The initial atlas $A_{t}^{0}$ is computed as the voxelwise median over all images $I_{i,t}$ in a time window $\tau\in[t-\delta,t+\delta]$. The median was chosen over the mean to produce a sharper initial atlas, as previously shown (Bastiaansen et al., 2022b). $\delta$ is a hyperparameter of the network.
@@ 2.3 Network architecture
For the nonrigid registration network, the default architecture of VoxelMorph was used (Balakrishnan et al., 2019). The network consists of an encoder, with convolutional layers with a stride of 2, a decoder with convolutional layers with a stride of 1, skip connections, and an upsampling layer, followed by two convolutional layers at full resolution with 16 filters. The encoder consists of one layer with 16 filters and is followed by 3 layers with 32 filters. The decoder consists of 5 layers with 32 filters.
To ensure invertibility of the deformation $\phi_{i,t}^{-1}$, we use diffeomorphic nonrigid deformations. Following Dalca et al. (2019), we use a stationary velocity field representation, meaning that $\phi_{{i,j}}$ is obtained by integrating the velocity field $\nu_{i,j}$ and $\phi_{i,j}^{-1}$ is obtained by integrating the negated velocity field $-\nu_{i,j}$ (Ashburner, 2007).
2.4 Optimization
The two networks were trained end-to-end using the ADAM optimizer with a learning rate of $10^{-6}$. The learning rate was chosen based on preliminary experiments. Training was conducted for 500 epochs with a batch size of one. During each epoch, every image in the training set was passed through the network once to update the model weights. After each epoch, the validation set was processed to compute the validation loss, which was used to monitor convergence and evaluate performance.
@@ 2.6 Implementation details
The network was trained on a RTX 2080 Ti 11 GB GPU with an AMD EPYC 7452 CPU using 14 GB of RAM. The code is publicly available at: https://gitlab.com/radiology/prenatal-image-analysis/4D_human_embryonic_brain_atlas
@@ 2.7 Voxel-based morphometry (VBM)
To perform VBM in a 4D reference space, the statistical comparison between subgroups can be performed at each gestational day. However, since we have limited data at some gestational days (less than 10 images), we choose to perform the statistical comparison within non-overlapping time windows of $12\delta+$ 1 days, for the same value of $\delta$ as used to create the initial atlas. Alternatively, the statistical comparison could be performed over the whole time period at once. We choose to perform the analysis per time window since the spatial correspondence throughout weeks 8 to 12 is limited by the appearance of structures (e.g., choroid plexus early week 9) and large changes in the curvature of the embryonic body. Within smaller time windows the spatial correspondence is maintained due to the time-dependent voxel sizes used and Gaussian smoothing of the Jacobian determinant.
@@ 4.1 Evaluation metrics
Registration accuracy was evaluated by calculating the Dice Similarity Coefficient (DSC) score between $S_{I_{i,t}}$ and $S_{A_{t}}\circ\phi_{i,t}^{-1}$ (Dice, 1945). The segmentation of the atlas $S_{A_{t}}$ was obtained by including all voxels in $A_{t}$ with image intensity $\geq 0.1$, followed by a binary opening and closing to remove noise and include the hypo-intense brain ventricles. Both the opening and closing were done using a ball with a radius of 10 voxels. The regularity of $\phi_{i,t}$ was evaluated by calculating the percentage of voxels having a non-positive Jacobian determinant %$\%|J|\leq 0$ within the mask $S_{A_{t}}$. For both the DSC and %$\%|J|\leq 0$ the mean and standard deviation over the validation and/or test set were reported.
@@
Finally, differences between generated atlases were evaluated by calculating the structural similarity index measure (SSIM) (Wang et al., 2004). The SSIM ranges from -1 to 1, where a value of 1 indicates perfect similarity, and 0 indicates no similarity. The SSIM is reported as $\text{SSIM}(\bar{A},A_{t})$, with $\bar{A}$ the reference atlas to which $A_{t}$ is compared. We reported the mean and standard deviation over the gestational days $t$ for $\text{HV}_{\text{error}}(A_{t},t)$, $\text{sharpness}(A_{t})$ and $\text{SSIM}(\bar{A},A_{t})$.
@@ 4.3 Experiment 2: Analysis best model
with $n(t)$ the number of images at gestational day $t$, and $u_{t}^{k}$ the $k^{\text{th}}$ displacement field at day $t$. We analyzed the mean and standard deviation over all voxels $x\in S_{A_{t}}$ of the average deformation $U_{t}(x)$. Secondly, to assess the learned generated deviation $A_{t}^{g}$, we visually inspected the shape and position of ventricles and brain structures in $A_{g}^{t}$. Thirdly, we evaluated the influence of atlas-generation dataset size. Dataset size was evaluated by splitting the atlas-generation set into $n=2,4,8$ subsets. When $n=2$ the subsets consist of approximately 100 subjects with 300 images, when $n=4$ of 75 subjects with 150 images, and when $n=8$ of 35 subjects with 80 images. We call the atlases $A_{t}^{n,m}$, with $n$ the number of subsets and $m=1,...,n$ the subset index. The $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ for every combination of $n\in\{2,4,8\}$ and $m=1,...,n$ was calculated. Here, $A_{t}$ is the atlas obtained using the complete atlas-generation set.
@@ 4.5 Experiment 4: Influence of maternal BMI on morphology
To assess whether the atlas can capture known differences in morphology, we studied the influence of maternal overweight and obesity during early pregnancy on first trimester brain development using VBM. We compared the following two populations: 1) subjects with a normal maternal BMI ($19\leq$ BMI <25)<25)< 25 ) and 2) subjects with a high maternal BMI (BMI $\geq 25$, overweight and obesity). Since there were only 7 subjects with a low BMI (BMI<19), we excluded those from this analysis. Additionally, for subjects who participated multiple times in the study, we retained only the first pregnancy to prevent including the same maternal environment twice. We tested whether there was a structural volumetric difference between the two populations by performing a t-test in a dataset of 819 ultrasound images of 392 subjects.
@@ 5.1 Experiment 1: Ablation and hyperparameters studies
Fourthly, evaluation the hyperparameter $\delta$, we observerd for $\delta\in\{1,2,3,4\}$, a lower $\text{HV}_{\text{error}}(A_{t},t)$, then using $\delta=0$. We observed that as $\delta$ increased, $\text{HV}_{\text{error}}(A_{t},t)$ decreased. For the $\text{sharpness}(A_{t})$, no clear pattern was found. We choose $\delta=3$, since this gave a good trade-off between the $\text{HV}_{\text{error}}(A_{t},t)$ and the $\text{sharpness}(A_{t})$. In Figure 4, we observed that for $\delta=3$, the atlas appears less noisy. In Figure 5, $\text{HV}_{\text{VR}}(t)$ and $\text{V}(A_{t})$ are shown. $\text{V}(A_{t})$ is strictly increasing, which indicates temporally consistent volumetric growth.
@@
Table 2: Results for the comparison of hyperparameters in Experiment 1. For %$\%|\nabla J|\leq 0$ and DSC the mean and standard deviation over the validation and/or test set are given. For $\text{SSIM}(A_{t}^{0},A_{t})$, $\text{HV}_{\text{error}}(A_{t},t)$ and $\text{sharpness}(A_{t})$ the mean and standard deviation over all gestational days $t$ is given.
@@
Figure 4: Atlas $A_{t}$ at different gestational days for different hyperparameter values tested in Experiment 1. All images are shown in the the mid-sagittal plane. The arrows indicate relevant visible brain structures. LV = lateral ventricle, 3V = third ventricle (cavity of the diencephalon), M = cavity of the mesencephalon, 4V = fourth ventricle (cavity of the rhombencephalon), CP = choroid plexus, CER = cerebellum.
@@ 5.2 Experiment 2: Analysis best model
Figure 8 gives the $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ for $n\in\{2,4,8\}$ and $m=1,...,n$. The error bars indicate the standard deviation over the gestational days for each atlas $A^{n,m}_{t}$. For both metrics, we observed that when $n$ gets bigger the metric is impacted negatively, and the spread over time gets wider. Furthermore, the variation among the $m$ generated atlases increased as $n$ increased. Figure 9 shows a visualization of the different atlases. Despite the lower SSIM as $n$ increased, in most cases, structures remained visible within the brain. The decrease in SSIM was mainly due to noise.
@@
Figure 8: Evaluation of the influence of atlas-generation dataset size in Experiment 2. Shown are the $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ with the number of subsets $n\in\{2,4,8\}$ and subset index $m=1,...,n$.

### v2

@@ Abstract
Early brain development is crucial for lifelong neurodevelopmental health. However, current clinical practice offers limited knowledge of normal embryonic brain anatomy on ultrasound, despite the brain undergoing rapid changes within the time-span of days. To provide detailed insights into normal brain development and identify deviations, we created the 4D Human Embryonic Brain Atlas using a deep learning-based approach for groupwise registration and spatiotemporal atlas generation. Our method introduced a time-dependent initial atlas and penalized deviations from it, ensuring age-specific anatomy was maintained throughout rapid development. The atlas was generated and validated using 831 3D ultrasound images from 402 subjects in the Rotterdam Periconceptional Cohort, acquired between gestational weeks 8 and 12. We evaluated the effectiveness of our approach with an ablation study, which demonstrated that incorporating a time-dependent initial atlas and penalization produced anatomically accurate results. In contrast, omitting these adaptations led to an anatomically incorrect atlas. Visual comparisons with an existing ex-vivo embryo atlas further confirmed the anatomical accuracy of our atlas. In conclusion, the proposed method successfully captures the rapid anotomical development of the embryonic brain. The resulting 4D Human Embryonic Brain Atlas provides a unique insights into this crucial early life period and holds the potential for improving the detection, prevention, and treatment of prenatal neurodevelopmental disorders.
Keywords spatiotemporal atlas $\cdot$ embryonic brain $\cdot$ brain development $\cdot$ groupwise image registration $\cdot$ first trimester ultrasound $\cdot$ deep learning
@@ 1 Introduction
We extended the work by Dalca et al. (2019), which was evaluated for Alzheimer’s disease characterized by relatively subtle anatomical changes over time. We extend their approach by introducing a time-dependent initial atlas and penalizing deviations from it, ensuring that rapid anatomical changes occurring over a span of days in first trimester brain development are preserved. The main contributions of research presented in this paper are:
@@
    A deep learning–based approach for spatiotemporal atlas creation [^1] Preliminary results were published in Bastiaansen et al. (2022b). There, we built upon the work by Dalca et al. (2019), but instead of jointly registering images to and generating the atlas, the images were registered to the atlas and the atlas was updated iteratively. adapted to account for rapid anatomical changes, validated through an ablation study showing that incorporating a time-dependent initial atlas and penalization of deviations improve anatomical accuracy compared to models without these adaptations.
@@
    The 4D Human Embryonic Brain Atlas, the first spatiotemporal brain atlas of the first trimester based on ultrasound imaging, which provides a framework for analyzing the influence of parental and lifestyle factors on early brain development, supports early detection of congenital anomalies, and serves as a clinically valuable visualization tool to enhance parental understanding and engagement.
@@
    An evaluation of the anatomical correctness of the resulting atlas by comparing its volume to a reference head volume (HV) curve established in previous research (Koning et al., 2016), and through visual comparison to the 3D Embryo Atlas by de Bakker et al. (2016) and the Fetal Brain Atlas by Namburete et al. (2023).
@@
    A study assessing whether the atlas can capture known morphological differences associated with maternal overweight and obesity during early pregnancy using Voxel-Based Morphometry (Ashburner and Friston, 2000).
@@ 2.1 Atlas $A_{t}$ and images $I_{i,t}$
Figure 1: Overview of the proposed method and network architectures for spatiotemporal atlas generation. The atlas generator network takes as input the gestational age $t$ (in days, normalized between 0 and 1) and the initial atlas $A_{t}^{0}$, and outputs the generated atlas $A_{t}=A_{t}^{0}+A_{t}^{g}$. It consists of a dense layer followed by three upsampling convolutional layers with 32 filters, with the number of upsampling layers chosen such that the final resolution matches that of the input images $I_{i,t}$. A final convolutional layer at full resolution produces $A_{t}^{g}$. The nonrigid registration network follows the VoxelMorph architecture (Balakrishnan et al., 2019) and includes an encoder with one convolutional layer of 16 filters and three layers with 32 filters (stride 2), a decoder with five upsampling convolutional layers with 32 filters, skip connections, and two convolutional layers with 16 filters at full resolution. Diffeomorphic deformations are obtained using a stationary velocity field representation, where the deformation field $\phi_{i,t}$ is computed by integrating the velocity field $\nu_{i,t}$, and the inverse deformation $\phi_{i,t}^{-1}$ by integrating its negation $\nu_{i,t}$ (Ashburner, 2007).
@@ 2.2 The time-dependent initial atlas $A_{t}^{0}$
The network generates the atlas by learning the deviation $A_{t}^{g}$ from the initial atlas, which is an input of the network together with the time $t$. We create a time-dependent initial atlas, instead of using a single initial atlas as proposed by Dalca et al. (2019), to account for the rapid morphological changes (e.g., the appearance of the choroid plexus, change of curvature of the embryonic body). The initial atlas $A_{t}^{0}$ is computed as the voxelwise median over all images $I_{i,t}$ in a time window $\tau\in[t-\delta,t+\delta]$. The median was chosen over the mean to produce a sharper initial atlas, as previously shown (Bastiaansen et al., 2022b). $\delta$ is a hyperparameter of the network.
@@ 2.3 Network architecture
For the nonrigid registration network, the default architecture of VoxelMorph was used (Balakrishnan et al., 2019). The network consists of an encoder, with convolutional layers with a stride of 2, a decoder with upsampling convolutional layers, skip connections, followed by two convolutional layers at full resolution with 16 filters. The encoder consists of one layer with 16 filters and is followed by 3 layers with 32 filters. The decoder consists of 5 upsampling layers with 32 filters.
To ensure invertibility of the deformation $\phi_{i,t}^{-1}$, we use diffeomorphic nonrigid deformations. Following Dalca et al. (2019), we use a stationary velocity field representation, meaning that $\phi_{{i,j}}$ is obtained by integrating the velocity field $\nu_{i,j}$ and $\phi_{i,j}^{-1}$ is obtained by integrating the negated velocity field $\nu_{i,j}$ (Ashburner, 2007).
@@ 2.5 Implementation details
The two networks were trained end-to-end using the ADAM optimizer with a learning rate of $10^{-6}$, chosen based on preliminary experiments. Training proceeded for 500 epochs with a batch size of one, with each image in the training set processed once per epoch to update the model weights. Validation loss was computed after each epoch to monitor convergence and evaluate performance. The training was carried out on an RTX 2080 Ti 11 GB GPU with an AMD EPYC 7452 CPU and 14 GB of RAM. The code used for these experiments is publicly available at https://gitlab.com/radiology/prenatal-image-analysis/4D_human_embryonic_brain_atlas .
@@ 2.6 Voxel-based morphometry (VBM)
To perform VBM in a 4D reference space, the statistical comparison between subgroups can be performed at each gestational day. However, since we have limited data at some gestational days (less than 10 images), we choose to perform the statistical comparison within non-overlapping time windows of $2\delta+1$ days, for the same value of $\delta$ as used to create the initial atlas. Alternatively, the statistical comparison could be performed over the whole time period at once. We choose to perform the analysis per time window since the spatial correspondence throughout weeks 8 to 12 is limited by the appearance of structures (e.g., choroid plexus early week 9) and large changes in the curvature of the embryonic body. Within smaller time windows the spatial correspondence is maintained due to the time-dependent voxel sizes used and Gaussian smoothing of the Jacobian determinant.
@@ 4.1 Evaluation metrics
Registration accuracy was evaluated by calculating the Dice Similarity Coefficient (DSC) score between $S_{I_{i,t}}$ and $S_{A_{t}}\circ\phi_{i,t}^{-1}$ (Dice, 1945). The segmentation of the atlas $S_{A_{t}}$ was obtained by including all voxels in $A_{t}$ with image intensity $\geq 0.1$, followed by a binary opening and closing to remove noise and include the hypo-intense brain ventricles. Both the opening and closing were done using a ball with a radius of 10 voxels. The regularity of $\phi_{i,t}$ was evaluated by calculating the percentage of voxels having a non-positive Jacobian determinant $\%|J|\leq 0$ within the mask $S_{A_{t}}$. For both the DSC and $\%|J|\leq 0$ the mean and standard deviation over the validation and/or test set were reported.
@@
Finally, differences between generated atlases were evaluated by calculating the structural similarity index measure (SSIM) (Wang et al., 2004). The SSIM ranges from -1 to 1, where a value of 1 indicates perfect similarity, and 0 indicates no similarity. The SSIM is reported as $\text{SSIM}(\bar{A},A_{t})$, with $\bar{A}$ the reference atlas to which $A_{t}$ is compared.
We reported the mean and standard deviation over the gestational days $t$ for $\text{HV}_{\text{error}}(A_{t},t)$, $\text{sharpness}(A_{t})$ and $\text{SSIM}(\bar{A},A_{t})$.
@@ 4.3 Experiment 2: Analysis best model
with $n(t)$ the number of images at gestational day $t$, and $u_{t}^{k}$ the $k^{\text{th}}$ displacement field at day $t$. We analyzed the mean and standard deviation over all voxels $x\in S_{A_{t}}$ of the average deformation $U_{t}(x)$. Secondly, to assess the learned generated deviation $A_{t}^{g}$, we visually inspected the shape and position of ventricles and brain structures in $A_{g}^{t}$. Thirdly, we evaluated the influence of atlas-generation dataset size. Dataset size was evaluated by splitting the atlas-generation set into $n=2,4,8$ subsets. When $n=2$ the subsets consist of approximately 100 subjects with 300 images, when $n=4$ of 75 subjects with 150 images, and when $n=8$ of 35 subjects with 80 images. We call the atlases $A_{t}^{n,m}$, with $n$ the number of subsets and m=1,,nm=1,...,n the subset index. The $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ for every combination of $n\in\{2,4,8\}$ and m=1,,nm=1,...,n was calculated. Here, $A_{t}$ is the atlas obtained using the complete atlas-generation set.
@@ 4.5 Experiment 4: Influence of maternal BMI on morphology
To assess whether the atlas can capture known differences in morphology, we studied the influence of maternal overweight and obesity during early pregnancy on first trimester brain development using VBM. We compared the following two populations: 1) subjects with a normal maternal BMI ($19\leq$ BMI <25)<25) and 2) subjects with a high maternal BMI (BMI $\geq 25$, overweight and obesity). Since there were only 7 subjects with a low BMI (BMI<19), we excluded those from this analysis. Additionally, for subjects who participated multiple times in the study, we retained only the first pregnancy to prevent including the same maternal environment twice. We tested whether there was a structural volumetric difference between the two populations by performing a t-test. Because maternal overweight and obesity may differently influence brain development, we additionally performed subgroup analyses, comparing normal BMI versus overweight ($25\leq$ BMI <30<30), normal BMI versus obese (BMI $\geq 30$), and overweight versus obese.
@@ 5.1 Experiment 1: Ablation and hyperparameters studies
Fourthly, we evaluated the hyperparameter $\delta\in{1,2,3,4}$. Increasing $\delta$ reduced the $\text{HV}_{\text{error}}(A_{t},t)$, while no clear trend was observed for $\text{sharpness}(A_{t})$. However, sharpness was slightly higher for $\delta=3$ than for $\delta=4$. Visual inspection revealed that for $\delta=4$, the choroid plexus appeared as early as 56 days, similar to $\delta=\infty$, which is anatomically incorrect (see Supplementary material 2 Figure 14). We therefore selected $\delta=3$, as it provided a good balance between volumetric accuracy, image sharpness, and anatomical plausibility. As shown in Figure 4, the atlas at $\delta=3$ appears less noisy, and Figure 5 demonstrates that $\text{V}(A_{t})$ increases monotonically, indicating temporally consistent volumetric growth.
@@
Table 2: Results for the comparison of hyperparameters in Experiment 1. For $\%|\nabla J|\leq 0$ and DSC the mean and standard deviation over the validation and/or test set are given. For $\text{SSIM}(A_{t}^{0},A_{t})$, $\text{HV}_{\text{error}}(A_{t},t)$ and $\text{sharpness}(A_{t})$ the mean and standard deviation over all gestational days $t$ is given.
@@
Figure 4: Atlas $A_{t}$ at different gestational days for different hyperparameter values tested in Experiment 1. All images are shown in the mid-sagittal plane. The arrows indicate relevant visible brain structures. LV = lateral ventricle, 3V = third ventricle (cavity of the diencephalon), M = cavity of the mesencephalon, 4V = fourth ventricle (cavity of the rhombencephalon), CP = choroid plexus, CER = cerebellum.
@@ 5.2 Experiment 2: Analysis best model
Figure 8 gives the $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ for $n\in\{2,4,8\}$ and m=1,,nm=1,...,n. The error bars indicate the standard deviation over the gestational days for each atlas $A^{n,m}_{t}$. For both metrics, we observed that when $n$ gets bigger the metric is impacted negatively, and the spread over time gets wider. Furthermore, the variation among the $m$ generated atlases increased as $n$ increased. Figure 9 shows a visualization of the different atlases. Despite the lower SSIM as $n$ increased, in most cases, structures remained visible within the brain. The decrease in SSIM was mainly due to noise.
@@
Figure 8: Evaluation of the influence of atlas-generation dataset size in Experiment 2. Shown are the $\text{SSIM}(A_{t},A_{t}^{n,m})$ and $\text{HV}_{\text{error}}(A_{t}^{n,m},t)$ with the number of subsets $n\in\{2,4,8\}$ and subset index m=1,,nm=1,...,n. Dots indicate the mean and error bars indicate the standard deviation.